23 y?z? sin(x4) dx dy dz
Answer #1


V 2. 2 3 2 3 ² 2 3 sin (27) sin (x7) dx dy dz 0 y-axis to x x = 2 c=4 x=y y=0 to y=2 -y=2 lin AJA X Xmaxid 5 1 2 2 3 4² 2 Sin an 2 3 x 2 Sin (x) dx de 23 3 VO 2 let 93 x t x 23 sin (x) dx dz 4x² dx = at 2010 x 3 dx = 4 - dt 4 2 T 23 sint at de 200 t=023 T 48 28 lost at let 2 - 12 4 t z d2= dt 3 TI 23 230 48 sint ! +1 98 = 230 23 (0-0) 48 48 - 231 Ans 98